Draw 25 samples from a distribution of IQ scores that are normally distributed with a mean of 100 and standard deviation of 15. What is the probability that the mean of 25 randomly drawn IQ scores will exceed 103 points?


1
Expert's answer
2021-11-29T16:27:50-0500

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{103-100}{15/\sqrt{25}}=1"


"P(\overline{x}>103)=P(z>1)=1-P(z<1)=1-0.8413=0.1587"